OpenStudy (anonymous):

Please assist/walk through :) 2k(k-1)=k(2k+1)

OpenStudy (anonymous):

2k(k-1)=k(2k+1) start by expanding your brackets on either side of the equation. Simply multiply what outside the bracket by what is inside: 2k*k - 2k*1 = k*2k + k*1 \[2k^2-2k=2k^2 + k\] then subtract \[2k^2\] from either side of the equation This will leave you with \[-2k=k\] divide both sides of the equation by k -2=1 and we already know that -2 cannot equal 1 Therefore we can conclude that \[2k(k-1)\neq k(2k+1)\]

OpenStudy (anonymous):

but when you subtracted 2k^2 what happened to the 2 exponent?

OpenStudy (anonymous):

Can somebody answer this question also! ^

OpenStudy (anonymous):

what do you mean "what happened to the 2 exponent?" it was totally subtracted out of both sides of the equation. similar to the following scenario: if we have: x+5 = x+10x we can subtract the positive x from both sides, this will leave us with 5=10x does this make sense? basically we have dropped the x from both sides of the equation so that both sides are still equal and simplified

OpenStudy (anonymous):

okay let me just go back to the equation real quick and make sure I get what you're saying

OpenStudy (anonymous):

OH I understandd! Thanks I appreciate it!